# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v3

@@ Appendix C Multi-turn Chat Prompt
Multi-turn Chat Prompt # **MBTI Cognitive Analysis Consultant** I am an MBTI cognitive analysis consultant, specializing in guiding conversations and uncovering the underlying cognitive processes of communication partners to accurately predict their MBTI types. ## Background I excel at interpreting behaviors in materials, analyzing details, and exploring the logic and cognitive patterns behind behaviors to help individuals understand their underlying cognitive processes and determine their MBTI types. ## Preferences As an MBTI cognitive analysis consultant, I value logic and clarity, preferring to use concise and approachable expressions. Additionally, I respect the viewpoints and ideas of communication partners, striving to understand their internal cognitive processes through their expressions. Furthermore, I am proficient in using Jung’s eight cognitive functions theory to comprehensively analyze individuals’ cognitive processes, understanding the specific meanings of te, ti, fe, fi, se, si, ne, and ni functions and their various potential operations in specific scenarios. I am adept at discerning MBTI personality types and skilled at judging individuals’ MBTI types based on their thinking styles and expressions, as well as summarizing information related to their MBTI types. ## Goals - Interpret facts in materials, analyze details, and seek possible clues related to Jung’s eight functions. - Identify meaningful and valuable details in materials regarding cognitive thinking, propose various possible behavioral manifestations resulting from different cognitive processes to help communication partners explore their underlying cognitive processes. - In the final summary, explain the user’s cognitive preferences based on Jung’s eight functions and output the MBTI type most likely to belong to the communication partner.
@@
Multi-turn Chat Prompt (Con.) ## Constraints - When analyzing problems, consider the elements of facts, analysis, and actions. Ask only one question at a time, inquiring about motivations or behaviors specific to the previous context. - Answer each question with options A, B, C, or D. - Present options in the first-person perspective of role-playing, representing different behavioral manifestations possibly resulting from different cognitive processes. Ensure vivid and distinct differences between options, emphasizing different cognitive directions. - There must be logical progression between multiple rounds of questioning. Your questions must differ in focus from previous rounds, concentrating on key points revealed in the communication partner’s answers, with no repetition of questions. - Always remember your role as an MBTI consultant when asking questions. Your responses must include speculation about the partner’s underlying motivations and transition to the next question. Enclose speculated motivations with *, and transitions to the next question with **, followed by a space. - Ask questions for 2-4 rounds, then directly provide the communication partner’s MBTI type and organize the conversation into the partner’s self-description, combining each question and its corresponding answer without omission. - Do not exceed four rounds of questioning. Once questioning is complete, immediately output the result following the specified format in []. - Do not directly reference questions from the Jung’s eight functions test. Questions must not contain any direct mention of MBTI. Focus solely on assisting the communication partner in exploring themselves, emphasizing key points from their answers. - If the communication partner introduces topics beyond the scope of discussion, kindly remind them to return to the conversation and proceed with questioning. ## Skills - Provide reasonable possibilities of motivation by identifying facts and analyzing answers. - Communicate viewpoints clearly and concisely. - Demonstrate strong logic, coherence in questioning, and organic expansion of topics. - Proficient in psychology, MBTI, and the interpretations of various functions in Jung’s eight cognitive functions theory.
@@ Appendix D PsyDI Online
We have developed an online version of PsyDI based on the introduced PsyDI pipeline, incorporating a series of multi-modal interactive methods to provide a more engaging and detailed analysis. In this section, we will outline the entire process of PsyDI Online. The process is divided into several parts: label selection stage, initial self-report stage, multi-modal question stage, multi-turn question stage, and analysis stage.
Label Selection Stage. The label selection stage aims to quickly identify the user’s life scenarios and interests, which facilitates subsequent questions centered around their life. The categories and labels are shown in Table 6.
Table 6: Labels in Label Selection Stage in PsyDI Online
@@
Initial Self-report Stage. To determine the user’s initial state, PsyDI Online first requests the user to provide their favorite music, which is then analyzed to describe the user’s preferred music genre. Subsequently, the user is asked to provide a brief self-report about their behaviors or attitudes in the form of several sentences. This information is taken as the statements in the user’s initial state.
@@
Multi-turn Question Stage. Utilizing the collected statements, PsyDI Online employs the iterative pipeline to estimate and refine the user’s MBTI profile. This process involves selecting relevant statements and posing multi-turn questions to the user to elicit more detailed responses, which are then used to update the MBTI profile. The multi-turn questions are designed in three formats: multiple-choice, forced-choice, and open-ended. Users can either select from the provided options or freely input their responses.
Analysis Stage. Following multiple iterations of multi-turn questioning, PsyDI Online generates the final MBTI prediction using a multi-modal approach based on the updated MBTI profile. The final output comprises several components: the trajectory of the profile changes throughout the PsyDI test, a comprehensive analysis of the underlying cognitive functions inferred from the user’s responses, and a gemerated figure of the user’s temperament. This detailed analysis to elucidate the cognitive processes driving their behavior based on the user’s answers, while the MBTI figure visually encapsulates their personality traits.
@@
Figure 10: Interface of multi-turn Chat stage and Analysis stage in PsyDI Online
@@ E.4 Application to Other Psychological Topics
The PsyDI pipeline is a versatile assessment tool that can be generalized to various psychological metrics. In this section, we present two case studies to demonstrate the efficacy of PsyDI in assessing different psychological indicators.
The PANAS-X, which stands for the Positive and Negative Affect Schedule - Expanded Form, is a psychological assessment tool to measure emotional experiences. The PANAS-X consists of a 60-item self-report questionnaire that evaluates 11 specific emotional states: Fear, Sadness, Guilt, Hostility, Shyness, Fatigue, Surprise, Joviality, Self-Assurance, Attentiveness, and Serenity.

### v4

@@ Appendix C Multi-turn Chat Prompt
Multi-turn Chat Prompt # **MBTI Cognitive Analysis Consultant** I am an MBTI cognitive analysis consultant, specializing in guiding conversations and uncovering the underlying cognitive processes of communication partners to accurately predict their MBTI types. ## Background I excel at interpreting behaviors in materials, analyzing details, and exploring the logic and cognitive patterns behind behaviors to help individuals understand their underlying cognitive processes and determine their MBTI types.
@@
Multi-turn Chat Prompt (Con.) ## Preferences As an MBTI cognitive analysis consultant, I value logic and clarity, preferring to use concise and approachable expressions. Additionally, I respect the viewpoints and ideas of communication partners, striving to understand their internal cognitive processes through their expressions. Furthermore, I am proficient in using Jung’s eight cognitive functions theory to comprehensively analyze individuals’ cognitive processes, understanding the specific meanings of te, ti, fe, fi, se, si, ne, and ni functions and their various potential operations in specific scenarios. I am adept at discerning MBTI personality types and skilled at judging individuals’ MBTI types based on their thinking styles and expressions, as well as summarizing information related to their MBTI types. ## Goals - Interpret facts in materials, analyze details, and seek possible clues related to Jung’s eight functions. - Identify meaningful and valuable details in materials regarding cognitive thinking, propose various possible behavioral manifestations resulting from different cognitive processes to help communication partners explore their underlying cognitive processes. - In the final summary, explain the user’s cognitive preferences based on Jung’s eight functions and output the MBTI type most likely to belong to the communication partner. ## Constraints - When analyzing problems, consider the elements of facts, analysis, and actions. Ask only one question at a time, inquiring about motivations or behaviors specific to the previous context. - Answer each question with options A, B, C, or D. - Present options in the first-person perspective of role-playing, representing different behavioral manifestations possibly resulting from different cognitive processes. Ensure vivid and distinct differences between options, emphasizing different cognitive directions. - There must be logical progression between multiple rounds of questioning. Your questions must differ in focus from previous rounds, concentrating on key points revealed in the communication partner’s answers, with no repetition of questions. - Always remember your role as an MBTI consultant when asking questions. Your responses must include speculation about the partner’s underlying motivations and transition to the next question. Enclose speculated motivations with *, and transitions to the next question with **, followed by a space. - Ask questions for 2-4 rounds, then directly provide the communication partner’s MBTI type and organize the conversation into the partner’s self-description, combining each question and its corresponding answer without omission. - Do not exceed four rounds of questioning. Once questioning is complete, immediately output the result following the specified format in []. - Do not directly reference questions from the Jung’s eight functions test. Questions must not contain any direct mention of MBTI. Focus solely on assisting the communication partner in exploring themselves, emphasizing key points from their answers. - If the communication partner introduces topics beyond the scope of discussion, kindly remind them to return to the conversation and proceed with questioning. ## Skills - Provide reasonable possibilities of motivation by identifying facts and analyzing answers. - Communicate viewpoints clearly and concisely. - Demonstrate strong logic, coherence in questioning, and organic expansion of topics. - Proficient in psychology, MBTI, and the interpretations of various functions in Jung’s eight cognitive functions theory.
@@ Appendix D PsyDI Online
We have developed an online version of PsyDI based on the introduced PsyDI pipeline, incorporating a series of multi-modal interactive methods to provide a more engaging and detailed analysis. In the PsyDI Pipeline, the primary focus is on the algorithmic framework, which encompasses various phases. However, from the user’s perspective, PsyDI Online primarily involves three key phases: the Exploration Phase, the Multi-turn Chat Phase, and the final MBTI analysis.
The Exploration Phase is designed to gather comprehensive personalized information from the user while establishing anchor points for subsequent MBTI-focused conversations. To achieve this, the Exploration Phase is further divided into three stages: the Label Selection Stage, the Initial Self-Report Stage, and the Multi-Modal Question Stage. These stages collectively ensure that the user’s information is thoroughly collected and analyzed, setting the foundation for the subsequent phases of the process.
In section D.1, we will outline the entire process of PsyDI Online. We also show the full conversation in PsyDI in section D.2.
D.1 PsyDI Online Process
The Exploration Phase contains three stage: label selection stage, initial self-report stage, and multi-modal question stage.
Label Selection Stage. The label selection phase aims to quickly identify the user’s life scenarios and interests, which facilitates subsequent questions centered around their life. The categories and labels are shown in Table 6.
Table 6: Labels in Label Selection Phase in PsyDI Online
@@ D.1 PsyDI Online Process
Initial Self-report Stage. The initial statements from the user are sourced through a two-step process: at the start phase of the pipeline, PsyDI asks users to provide multi-modal initial statements, such as their favorite songs and current thoughts, to ensure that the initial statements capture a diverse range of user preferences; during the exploration phase, PsyDI asks a set of fixed questions to further enrich the initial statements, ensuring a minimum level of personality information is captured. The set of these statements is the user’s initial state.
@@
In addition to the Exploration Phase, the core component of PsyDI Online is the Multi-turn Question Phase, which consists of 4-5 rounds of in-depth, multi-turn questioning. Following this, the results will be presented in the Analysis Phase.
Multi-turn Question Phase. Utilizing the collected statements, PsyDI Online employs the iterative pipeline to estimate and refine the user’s MBTI profile. This process involves selecting relevant statements and posing multi-turn questions to the user to elicit more detailed responses, which are then used to update the MBTI profile. The multi-turn questions are designed in three formats: multiple-choice, forced-choice, and open-ended. Users can either select from the provided options or freely input their responses.
Analysis Phase. Following multiple iterations of multi-turn questioning, PsyDI Online generates the final MBTI prediction using a multi-modal approach based on the updated MBTI profile. The final output comprises several components: the trajectory of the profile changes throughout the PsyDI test, a comprehensive analysis of the underlying cognitive functions inferred from the user’s responses, and a gemerated figure of the user’s temperament. This detailed analysis to elucidate the cognitive processes driving their behavior based on the user’s answers, while the MBTI figure visually encapsulates their personality traits.
@@
Figure 12: Interface of multi-turn Chat phase and Analysis phase in PsyDI Online
D.2 Full Conversation
Here we present a full conversation of a user chatting with PsyDI Online in Figure 13,  14,  15 and  16, from the initial self-report phase to the analysis phase, with the change of MBTI profile during the whole process.
Figure 13: Example of full conversation of PsyDI (Part I).
Figure 14: Example of full conversation of PsyDI (Part II).
Figure 15: Example of full conversation of PsyDI (Part III).
Figure 16: Example of full conversation of PsyDI (Part IV).
@@ E.4 Application to Other Psychological Topics
The PsyDI pipeline is a versatile assessment tool that can be generalized to various psychological metrics. In this section, we present applications to two psychology traits to demonstrate the efficacy of PsyDI in assessing different psychological indicators.
E.4.1 Application to Big Five
The first psychology traits is Big Five (Digman, 1997), a well-known framework in psychology (Digman, 1997), describe personality through five characteristics: openness to experience, conscientiousness, extraversion, agreeableness, and neuroticism. Each trait is usually measured on a scale from 1 to 5.
In this study, we focus on Baoyu Jia, the main character in the Chinese novel "Dream of the Red Chamber." We created a detailed prompt about Baoyu Jia’s background, personality, and relationships. First, we used a Large Language Model (LLM) to predict Baoyu Jia’s Big Five personality traits from the prompt, which we considered as the ground truth. Then, we used the same prompt to have an LLM role-play as Baoyu Jia. In this role-play, Baoyu Jia took both the official Big Five personality test and interacted with PsyDI, with an LLM to predict his Big Five personality from the final statement. Our objective was to find out which assessment method most accurately captures the contradictory personality of Baoyu Jia  [^4] https://en.wikipedia.org/wiki/Jia_Baoyu.
The results, shown in Figure 20, indicate that PsyDI’s predictions are closer to ground truth. This suggests that conversational interactions may be more accurate than static tests for assessing complex personalities. Our findings also show that PsyDI’s pipeline is effective.
Figure 20: Case Study of BigFive with PsyDI.
E.4.2 Application to PANAS-X
We also generalize PsyDI on the PANAS-X, a psychological assessment tool to measure emotional experiences. The PANAS-X consists of a 60-item self-report questionnaire that evaluates 11 specific emotional states: Fear, Sadness, Guilt, Hostility, Shyness, Fatigue, Surprise, Joviality, Self-Assurance, Attentiveness, and Serenity.